# Nesterov acceleration in benignly nonconvex landscapes

### v1

@@ 1. Introduction
On the other hand, strictly negative (but small) eigenvalues of the Hessian of the loss function have been observed close to the set of global minimizers experimentally by Sagun et al. (2017, 2018); Alain et al. (2018) and their presence has been explained theoretically by Wojtowytsch (2023). This poses questions about the use of momentum-based optimizers such as SGD with (heavy ball or Nesterov) momentum or Adam in the training of deep neural networks.
In overparametrized learning, the set of minimizers is a submanifold of high dimension in a usually much higher-dimensional space, leading to novel geometric phenomena compared to classical convex optimizaton.
Heuristically, as long as the objective function is convex in the direction towards the set of minimizers, small negative eigenvalues in directions parallel to the set of minimizers can safely be ignored: Tangential drift along the set of minimizers does not affect the decay of the objective function significantly. We prove that this intuition indeed applies in a continuous time model for gradient descent with momentum (Theorems 7 and 8) and for Nesterov’s time-stepping scheme (Theorems 11, 13 and 14) in deterministic optimization and stochastic optimization with bounded noise. With ‘multiplicative (state-dependent) noise’ motivated by overparametrized deep learning, we prove an analogous statement for a modified version of Nesterov’s algorithm (Theorem 15).
@@
In the remainder of this section, we briefly review how our work fits into the literature. In Section 2, we precisely state the assumptions under which we prove convergence at an accelerated rate and demonstrate how our work connects to optimization in deep learning. Our main results are presented in Section 3, both in discrete and continuous time. We discuss the implications and further directions in Section 4. Some technical details are postponed to the Appendix.
1.1. Previous work
@@
Gradient-based optimization was first proposed by Cauchy et al. (1847) in form of the gradient descent algorithm. Over a century later, momentum-based ‘accelerated’ algorithms were introduced by Hestenes & Stiefel (1952) for convex quadratic functions and by Nesterov (1983) for general smooth and convex objective functions. Nesterov’s work was generalized to non-smooth convex optimization by Beck & Teboulle (2009) and to stochastic smooth convex optimization among others by Nemirovski et al. (2009); Shamir & Zhang (2013); Jain et al. (2019); Laborde & Oberman (2020) for additive noise and by Liu & Belkin (2018); Even et al. (2021); Vaswani et al. (2019); Gupta et al. (2023) for multiplicatively scaling noise. See also (Ghadimi & Lan, 2012, 2013) for more information on accelerated stochastic gradient methods.
@@
Necoara et al. (2019); Hinder et al. (2020); Aujol et al. (2022); Fu et al. (2023); Wang & Wibisono (2023); Guminov et al. (2023) consider acceleration under the weaker condition that the objective function is $\gamma$-quasar or $(\gamma,\mu)$-strongly quasar-convex, i.e. the inequality
@@
holds for any $x\in\mathbb{R}^{d}$ and any minimizer $x^{*}$ of $f$. Compared to (strong) convexity, it relaxes the condition in two ways: It only considers pairs $(x,x^{*})$ rather than general pairs of points $x,y\in\mathbb{R}^{d}$, and it introduces a factor $\gamma$ into the inequality which may be strictly smaller than one. Still, it has geometric implications which may be too strong in the context of deep learning: In the strongly quasar-convex case, minimizers are unique, and in the quasar-convex case, sub-level sets are star-shaped with respect to any minimizer $x^{*}$ since $f(tx+(1-t)x^{*})$ is monotone increasing on $[0,1]$. Weaker, but still restrictive conditions such of quadratic growth type were considered in particular by Necoara et al. (2019).
In general deep learning applications, the set of minimizers of the loss function is a (generally curved) manifold, and tangential motion to the manifold can have important implications on the implicit bias of an algorithm (Li et al., 2021; Damian et al., 2021). Any notion that takes into account all minimizers is quite rigid for such tasks. A more realistic assumption is the ‘aiming condition’ of Liu et al. (2024) that
@@
Under the (comparatively weak) Polyak-Lojasiewicz (PL) condition, Yue et al. (2023) show that it is not possible to obtain an accelerated rate of convergence. A slower linear rate of convergence is established by Apidopoulos et al. (2022) in continuous time under the assumption that the objective function $f$ satisfies has an $L$-Lipschitz continuous gradient and satisfies the PL-inequality $2\mu\big{(}f-\inf f\big{)}\leq\|\nabla f\|^{2}$. The rate of convergence is
A stable time-discretization can generally be attained with effective step-size $1/\sqrt{L}$ for momentum methods as we explain below, suggesting that this corresponds to convergence at the non-accelerated linear rate $(1-\mu/2L)^{k}$ in discrete time. To the best of our knowledge, a rigorous proof is not yet available.
Our study can be seen as combining geometric ideas pertaining to $(1,\mu)$-quasar convexity and the aiming condition. To the best of our knowledge, it is the first study which takes into account tangential motion along the set of minimizers and obtains an accelerated rate of convergence.
2. Setting
2.1. Assumptions
@@
  4. (4)
    $f$ is first order $\mu$-strongly-convex with respect to the closest minimizer in $\mathcal{U}_{\alpha}$, i.e.
    (1) $\nabla f(x)\cdot\big{(}x-\pi(x)\big{)}\geq f(x)-f\big{(}\pi(x)\big{)}+\frac{\mu}{2}\,\|x-\pi(x)\|^{2}\qquad\forall\ x\in\mathcal{U}_{\alpha}.$
@@
Lemma 1.
Let $\mathcal{M}$ be a $C^{2}$-submanifold of $\mathbb{R}^{d}$ and $\mathcal{U}$ an open set containing $\mathcal{M}$ such that there exists a unique closest point projection $\pi:\mathcal{U}\to\mathcal{M}$. Let $x:(-\varepsilon,\varepsilon)\to\mathcal{U}$ a $C^{1}$-curve and $z(t):=\pi\circ x(t)$. Then $\langle\dot{x},\dot{z}\rangle\geq 0$ on $(-\varepsilon,\varepsilon)$.
2.2. Simple Examples
@@
PL condition $\mu$-sc wrt minimizer $\mu$-strongly convex
@@
Evidently, the geometric conditions and associated constants are quite different if $R\gg 1$. See Figure 1 for an illustration of $f$. Further details for the example and a comparison to less common notions such as quasar-convexity are given in Appendix C.
Figure 1. We visualize $f$ from Example 2 in the top row and its derivative in the bottom row with $R=2$ and $\varepsilon=0.2$ (left), $\varepsilon=0.1$ (middle) and $\varepsilon=0.05$ (right). In the left plot, $f$ has many local minimizers as the derivative crosses 00 an infinite number of times. In the middle plot, $f$ satisfies the PL condition, but not the first order strong convexity condition with respect to the unique global minimizer. In the right plot, $f$ is strongly first-order convex with respect to the unique global minimizer (which implies the PL condition). In all plots, $f$ is non-convex since $f^{\prime}$ is non-monotone.
@@
$f:\mathbb{R}^{d}\to\mathbb{R},\qquad f(x)=\tilde{f}_{2}(x_{1},\dots,x_{k})\cdot\tilde{f}(x_{k+1},\dots,x_{d}).$
Then $f$ is $a\mu$-strongly convex with respect to the closest minimizer $\pi(x)=(x_{1},\dots,x_{k},0,\dots,0)$, but not strongly convex since the minimizer is non-unique. Similarly, if $A:\mathbb{R}^{k}\to\mathbb{R}^{(d-k)\times(d-k)}$ is a function which takes values in the set of symmetric matrices with eigenvalues larger than $\mu$, then
@@
is $\mu$-strongly convex with respect to the closest minimizer, but generally non-convex.
@@
for $x\in\mathcal{U}_{\varepsilon}$, i.e. $f$ is first order $\mu$-strongly convex with respect to the closest minimizer.
On the other hand, $f$ is not convex unless $\mathcal{M}$ is. Otherwise, take $x_{1},x_{2}\in\mathcal{M}$ and $t\in(0,1)$ such that $tx_{1}+(1-t)x_{2}\notin\mathcal{M}$. Then the map $t\mapsto d^{2}\big{(}tx_{1}+(1-t)x_{2}\big{)}$ attains a maximum inside the interval $(0,1)$, meaning that $d^{2}$ cannot be convex.
@@
Figure 2. Left: The dashed red line connects two minimizers of the function $f$. Along the line, $f$ must achieve an interior local maximum. At this point, the Hessian $D^{2}f$ cannot be positive definite. Middle, Right: Optimization trajectories for Nesterov’s method (top) and its associated energy curve (bottom). The selection of limit point may depend crucially on optimization parameters: In the middle plot, we take 800 steps with stepsize $10^{-2}$ while on the right, we take 8,000 steps with stepsize $10^{-3}$ from the same initial point. The decay of $f(x_{t})$ is similar for both trajectories, but the limit points on the manifold of minimizers are far apart. The objective function is $f(x,y)=(x^{2}/2+3y^{2}-1)^{2}$.
2.3. Application to Deep Learning
An important class of objective functions are those which combine the geometric features of Examples 3 and 4. Such functions can be seen as geometric prototypes for loss functions in deep learning. Namely, consider a parametrized function class $h:\mathbb{R}^{p}\times\mathbb{R}^{d}\to\mathbb{R}$ of weights $w\in\mathbb{R}^{p}$ and data $x\in\mathbb{R}^{d}$ (e.g. a neural network) and the mean squared error (MSE) loss function
(2) $L_{y}:\mathbb{R}^{p}\to[0,\infty),\qquad L_{y}(w)=\frac{1}{2n}\sum_{i=1}^{n}\big{(}h(w,x_{i})-y_{i}\big{)}^{2},\qquad y=(y_{1},\dots,y_{n})\in\mathbb{R}^{n}.$
@@ 2.3. Application to Deep Learning
has full rank $n$ at every $x\in\mathcal{M}_{y}$ (for almost every $y$). Thus, all $n$ eigenvalues in direction orthogonal to $\mathcal{M}$ are non-zero. We prove the following in Appendix A.
@@
Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ is $C^{2}$-smooth and that =$\mathcal{M}=\{x\in\mathbb{R}^{d}:f(x)=\inf f\}$ is a closed $k$-dimensional $C^{2}$-submanifold of $\mathbb{R}^{d}$ (i.e. compact and without boundary). If $D^{2}f(x)$ has rank $d-k$ everywhere on $\mathcal{M}$, then there exist $\mu,\alpha>0$ such that there exists a $C^{1}$-smooth closest point projection $\pi:U_{\alpha}\to\mathbb{R}$ with $U_{\alpha}=\{x:f(x)<\alpha\}$ as in Section 2.1 and (1) holds for $f$ and $\pi$.
@@
In particular, $f,\mathcal{M}$ meet all conditions in Section 2.1. Note that by Lemma 5, the loss function $f$ cannot be convex since a compact manifold cannot be perfectly straight everywhere.
@@
Wojtowytsch (2023, Theorem 2.6) shows that the assumption that $\mathcal{M}$ is compact is a simplification and generally does not apply in deep learning. Local versions of Lemma 6 could be proved with $\mu,\alpha$ which are positive functions on the manifold, but not necessarily bounded away from zero. Naturally, this suffices in all cases where we provably remain in a local neighborhood in the course of optimization. We eschew this greater generality for the sake of geometric clarity and commit the pervasive sin of optimization theory for deep learning: We make global assumptions which can only be guaranteed locally. For a further discussion of geometric conditions in optimization and deep learning, see also Appendix C.3.
We illustrate in Figure 3 that our assumptions are locally reasonable in deep learning. We trained a fully connected neural network with 10 layers of width 35 and tanh activation to fit labels $y_{i}$ at $100$ randomly generated datapoints $x_{i}$ in $\mathbb{R}^{12}$. The small dataset size allowed us to use the exact gradient and loss function rather than stochastic approximations for a better exploration of the loss landscape. As the direction to the closest minimizer generally cannot be found, we use the gradient as a proxy and investigate the convexity of the function $\phi(t)=L(w+tg)$ for a point $w$ which is very close to the set of global minimizers of a loss function $L$ as in (2) and the direction $g=\nabla L(w)/\|\nabla L(w)\|$.
The labels were generated by a teacher network of width 20 and depth 7 with randomly initialized parameters. The student networks were trained for 10,000 epochs using stochastic gradient descent with Nesterov momentum and learning rate $\eta=0.005$, momentum $\rho=0.99$. At the end of training, the training loss was between $10^{-12}$ and $10^{-9}$ for all five runs. Second derivatives are approximated by second order difference quotients $\phi(t)\approx\frac{\phi(t+h)-2\phi(t)+\phi(t-h)}{h^{2}}$ for $h=0.01$. Similarly, the strong convexity parameter with respect to the global minimizer is estimated by $2\frac{\phi^{\prime}(t)t-\phi(t)+\inf\phi}{t^{2}}$ where $\phi^{\prime}(t)$ is estimated by $\frac{\phi(t+h)-\phi(t-h)}{2h}$.
It should be noted that due to the inherent randomness of the experiment, a variety of geometric behaviors was observed. Generally, the loss function is convex in gradient direction in a neighborhood of a global minimizer, but the size of the neighborhood may vary, as well as the magnitude of second derivatives and the steepness of the objective function. In several experiments, there were examples of convex but not strongly convex behavior in gradient directions. In some experiments, training failed and the minimum value achieved along the line $w+tg/\|g\|$ was larger than zero. The gradient is an imperfect approximation of the direction towards the minimizer and rounding errors may occur as the gradient is very small so close to the set of minimizers.
@@
Figure 3. We investigate the convexity of the function $\phi(t)=L(w+tg)$ for a point $w$ which is very close to the set of global minimizers of a loss function $L$ as in (2) and the direction $g=\nabla L(w)/\|\nabla L(w)\|$. We plot $\phi(t)$ (left), the second derivative of $\phi$ (middle), and an estimate of the parameter $\mu$ of strong convexity with respenct to the minimizer (right) for $t\in[-1,1]$. Evidently, $\phi$ is strongly convex in a neighborhood of the set of global minimizers and the first order convexity inequality with respect to the minimizer along the line yields (in some cases significantly) larger constants than the parameter of strong convexity found by the second derivative. Different colors correspond to different runs with different random initialization.
3. Main Contributions
3.1. Optimization in continuous time

### v2

# Nesterov acceleration in benignly non-convex landscapes

@@ 1 Introduction
Recall that a $C^{2}$-smooth function $f$ is convex if and only if its Hessian, $D^{2}f$, is positive semi-definite, i.e. has only non-negative eigenvalues. Strictly negative (but small) eigenvalues of the Hessian of the loss function have been observed close to the set of global minimizers experimentally by Sagun et al. (2017; 2018); Alain et al. (2018) and their presence has been explained theoretically by Wojtowytsch (2023). This poses questions about the use of momentum-based optimizers such as SGD with (heavy ball or Nesterov) momentum or Adam in the training of deep neural networks. In this work, we show that acceleration can be guaranteed for Nesterov’s method under much weaker geometric assumptions than (strong) convexity, in particular for certain objective functions that have non-unique and non-isolated minimizers and whose Hessian may have negative eigenvalues up to a certain size.
@@
In the remainder of this section, we briefly review how our work fits into the literature. In Section 2, we precisely state the assumptions under which we prove convergence at an accelerated rate and discuss how our work connects to optimization in deep learning. Our main results are presented in Section 3, both in discrete and continuous time. Some technical details are postponed to the appendix.
1.1 Previous work
@@
Gradient-based optimization was first proposed by Cauchy et al. (1847) in form of the gradient descent algorithm. Over a century later, momentum-based ‘accelerated’ algorithms were introduced by Hestenes & Stiefel (1952) for convex quadratic functions and by Nesterov (1983) for general smooth and convex objective functions. Nesterov’s work was generalized to non-smooth convex optimization by Beck & Teboulle (2009) and to stochastic smooth convex optimization among others by Nemirovski et al. (2009); Shamir & Zhang (2013); Jain et al. (2019); Laborde & Oberman (2020) for additive noise and by Liu & Belkin (2018); Even et al. (2021); Vaswani et al. (2019); Gupta et al. (2024) for multiplicatively scaling noise. See also (Ghadimi & Lan, 2012; 2013) for more information on accelerated stochastic gradient methods. While the heavy ball method is used extensively in deep learning, Lessard et al. (2016); Goujaud et al. (2023) prove that it does not generally achieve accelerated convergence for smooth strongly convex functions and may even diverge – see however (Kassing & Weissmann, 2024) for positive results under stronger smoothness assumptions
Accelerated gradient methods have been studied e.g. by Josz et al. (2023) under much weaker regularity conditions and weaker geometric conditions than (strong) convexity, namely the Kurdyka-Lojasiewicz (KL) condition. Under those weaker assumptions, it is at best possible to prove convergence to a local minimizer at a non-acclerated rate: Under the (comparatively weak) Polyak-Lojasiewicz (PL) condition, a special case of the KL condition, Yue et al. (2023) show that it is not possible to obtain an accelerated rate of convergence. A slower linear rate of convergence is established by Apidopoulos et al. (2022) in continuous time under the assumption that the objective function $f$ satisfies has an $L$-Lipschitz continuous gradient and satisfies the PL-inequality $2\mu\big{(}f-\inf f\big{)}\leq\|\nabla f\|^{2}$. The rate of convergence is
A stable time-discretization can generally be attained with effective step-size $1/\sqrt{L}$ for momentum methods (see Su et al., 2016, Section 2), suggesting convergence at the non-accelerated linear rate $(1-\mu/2L)^{k}$ in discrete time. To the best of our knowledge, no proof has been given yet.
@@
There have been several efforts to find a reasonable relaxation of convexity for which accelerated convergence can still be achieved. Hinder et al. (2020); Fu et al. (2023); Wang & Wibisono (2023); Guminov et al. (2023) consider acceleration under the weaker condition that the objective function is $\gamma$-quasar or $(\gamma,\mu)$-strongly quasar-convex, i.e. the inequality
@@
holds for any $x\in\mathbb{R}^{d}$ and any minimizer $x^{*}$ of $f$. Compared to (strong) convexity, it relaxes the condition in two ways: It only considers pairs $(x,x^{*})$ rather than general pairs of points $x,y\in\mathbb{R}^{d}$, and it introduces a factor $\gamma$ into the inequality which may be strictly smaller than one. Still, it has geometric implications which may be too strong in the context of deep learning: In the strongly quasar-convex case, minimizers are unique, and in the quasar-convex case, sub-level sets are star-shaped with respect to any minimizer $x^{*}$ since $f(tx+(1-t)x^{*})$ is monotone increasing on $[0,1]$ (see Lemma 24, Appendix C).
Accelerated rates of convergence were obtained by Necoara et al. (2019) in discrete time and by Aujol et al. (2022) in continuous time under the assumption that the objective function is both convex and quasi-strongly convex, and that it has a unique minimizer. Their results are generalized by Hermant et al. (2024) who allow for non-smooth composite optimization and consider $\gamma\in(0,1]$ rather than just $\gamma=1$. Unlike the present work, Hermant et al. (2024) require the uniqueness of minimizers and only study deterministic optimization. Curiously, despite the difference in settings, they independently find the same lower bound on Hessian eigenvalues that we require (compare e.g. Theorem 11 and (Hermant et al., 2024, Theorem 2)). See also (Aujol et al., 2024, Table 1) for an overview of theoretical guarantees of acceleration without strong convexity.
In overparametrized deep learning, the set of minimizers of the loss function is a (generally curved) manifold, and tangential motion to the manifold can have important implications on the implicit bias of an algorithm (Li et al., 2021; Damian et al., 2021). Any notion that takes into account all minimizers is quite rigid for such tasks. A more realistic assumption is the ‘aiming condition’ of Liu et al. (2024) that
@@
1.2 Our contribution
Our study can be seen as combining geometric ideas pertaining to $(1,\mu)$-quasar convexity and the aiming condition. Our assumption in (1) is equivalent to quasi-strong convexity (Necoara et al., 2019), but notably we do not make any additional assumptions about the convexity of the objective function or uniqueness of minimizers unlike prior works. Since we focus on the closest minimizer at any point in the trajectory, this presents a unique challenge compared to analyzing the distance from a fixed minimizer. As the current iterate $x_{t}$ moves, its projection onto the set of minimizers, $\pi(x_{t})$, also moves. This requires a modification of the usual Lyapunov function to account for the movement of $\pi(x_{t})$ (or the tangential movement of $x_{t}$ parallel to the set of minimizers) as well as the ‘drift’ in directions where the objective function is negatively curved. To the best of our knowledge, this is the first study which takes into account this tangential movement and obtains accelerated convergence.
In overparametrized learning, the set of minimizers of a loss function is a submanifold of high dimension in a usually much higher-dimensional space. Unless the manifold of minimizers is a linear space, the Hessian of the loss function is geometrically required to have negative eigenvalues in any neighborhood of the set of a minimizer where the manifold is curved. Still, accelerated methods in first order optimization have been found to be highly successful in deep learning. A common heuristic has been that as long as the objective function is convex in the direction towards the set of minimizers, small negative eigenvalues in directions parallel to the set of minimizers can safely be ignored: Tangential drift along the set of minimizers should not affect the decay of the objective function significantly. We prove that this intuition indeed applies in a continuous time model for gradient descent with momentum (Theorems 6 and 8) and for Nesterov’s time-stepping scheme (Theorems 11, 13 and 14) in deterministic optimization and stochastic optimization with bounded noise. With ‘multiplicative (state-dependent) noise’ motivated by overparametrized deep learning, we prove an analogous statement for a modified version of Nesterov’s algorithm (Theorem 15).
2 Setting
2.1 Assumptions
@@
  4. 4.
    $f$ satisfies the $\mu$-strong aiming condition (with respect to the closest minimizer) in $\mathcal{U}_{\alpha}$, i.e.
    $\nabla f(x)\cdot\big{(}x-\pi(x)\big{)}\geq f(x)-f\big{(}\pi(x)\big{)}+\frac{\mu}{2}\,\|x-\pi(x)\|^{2}\qquad\forall\ x\in\mathcal{U}_{\alpha}.$ (1)
@@
2.2 Simple Examples
@@
PL condition $\mu$-strongly aiming $\mu$-strongly convex
@@
Evidently, the geometric conditions and associated constants are quite different if $R\gg 1$. See Figure 1 for an illustration of $f$. Further details for the example and a comparison to less common notions such as quasar-convexity are given in Appendix C.2. We note that the example exploits the fact that $f$ is $C^{1,1}$- but not $C^{2}$-smooth: For $C^{2}$-functions, Rebjock & Boumal (2023) prove that the PL condition locally implies strong aiming condition.
Figure 1: We visualize $f$ from Example 1 in the top row and its derivative in the bottom row with $R=2$ and $\varepsilon=0.2$ (left), $\varepsilon=0.1$ (middle) and $\varepsilon=0.05$ (right). Left: $f$ has many local minimizers as the derivative crosses 00 an infinite number of times. Middle: $f$ satisfies the PL condition, but not the strong aiming condition. Right: $f$ is strongly aiming (with respect to the unique global minimizer, which implies the PL condition). In all plots, $f$ is non-convex since $f^{\prime}$ is non-monotone.
@@
$f:\mathbb{R}^{d}\to\mathbb{R},\qquad f(x)=\tilde{f}_{2}(x_{1},\dots,x_{k})\cdot\tilde{f}_{1}(x_{k+1},\dots,x_{d}).$
Then $f$ is $a\mu$-strongly aiming $\pi(x)=(x_{1},\dots,x_{k},0,\dots,0)$, but not strongly convex since the minimizer is non-unique. Similarly, if $A:\mathbb{R}^{k}\to\mathbb{R}^{(d-k)\times(d-k)}$ is a function which takes values in the set of symmetric matrices with eigenvalues larger than $\mu$, then
@@
is $\mu$-strongly aiming, but generally non-convex.
@@
for $x\in\mathcal{U}_{\varepsilon}$, i.e. $f$ is $\mu$-strongly aiming. On the other hand, $f$ is not convex unless $\mathcal{M}$ is. Otherwise, take $x_{1},x_{2}\in\mathcal{M}$ and $t\in(0,1)$ such that $tx_{1}+(1-t)x_{2}\notin\mathcal{M}$. Then the map $t\mapsto d^{2}\big{(}tx_{1}+(1-t)x_{2}\big{)}$ attains a maximum inside the interval $(0,1)$, meaning that $d^{2}$ cannot be convex.
@@
Figure 2: Left: The dashed red line connects two minimizers of the function $f$. Along the line, $f$ must achieve an interior local maximum. At this point, the Hessian $D^{2}f$ cannot be positive definite. Middle, Right: Optimization trajectories for Nesterov’s method (top) and its associated energy curve (bottom). The selection of limit point may depend crucially on optimization parameters: In the middle plot, we take 800 steps with stepsize $10^{-2}$ while on the right, we take 8,000 steps with stepsize $10^{-3}$ from the same initial point. The decay of $f(x_{t})$ is similar for both trajectories, but the limit points on the manifold of minimizers are far apart. The objective function is $f(x,y)=(x^{2}/2+3y^{2}-1)^{2}$.
2.3 Connection to Deep Learning
An important class of objective functions are those which combine the geometric features of Examples 2 and 3. Such functions can be seen as geometric prototypes for loss functions in overparametrized regression problems, such as in deep learning. Namely, consider a parametrized function class $h:\mathbb{R}^{p}\times\mathbb{R}^{d}\to\mathbb{R}$ of weights $w\in\mathbb{R}^{p}$ and data $x\in\mathbb{R}^{d}$ (e.g. a neural network) and the mean squared error (MSE) loss function
$L_{y}:\mathbb{R}^{p}\to[0,\infty),\qquad L_{y}(w)=\frac{1}{2n}\sum_{i=1}^{n}\big{(}h(w,x_{i})-y_{i}\big{)}^{2},\qquad y=(y_{1},\dots,y_{n})\in\mathbb{R}^{n}.$ (2)
@@ 2.3 Connection to Deep Learning
has full rank $n$ at every $x\in\mathcal{M}_{y}$ (for almost every $y$). Thus, all $n$ eigenvalues in direction orthogonal to $\mathcal{M}$ are non-zero. We prove the following in Appendix A. The same connection has been made e.g. by Rebjock & Boumal (2023) for $C^{2}$-functions and overparametrized regression problems.
@@
Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ is $C^{2}$-smooth and that =$\mathcal{M}=\{x\in\mathbb{R}^{d}:f(x)=\inf f\}$ is a closed $k$-dimensional $C^{2}$-submanifold of $\mathbb{R}^{d}$ (i.e. compact and without boundary). If $D^{2}f(x)$ has rank $d-k$ everywhere on $\mathcal{M}$, then there exist $\mu,\alpha>0$ such that there exists a $C^{1}$-smooth closest point projection $\pi:U_{\alpha}\to\mathbb{R}$ with $U_{\alpha}=\{x:f(x)<\alpha\}$ and $f$ is $\mu$-strongly aiming with $\pi$.
@@
In particular, $f,\mathcal{M}$ meet all conditions in Section 2.1. Note that by Lemma 4, the loss function $f$ cannot be convex since a compact manifold cannot be perfectly straight everywhere. Lemma 5 does not apply to networks with the non-smooth ReLU activation $\sigma(z)=\max\{z,0\}$, also here minimizers cannot be isolated due to the continuous scaling symmetry $\sigma(z)=\lambda^{-1}\sigma(\lambda z)$ for $\lambda>0$.
@@
Wojtowytsch (2023, Theorem 2.6) shows that the assumption that $\mathcal{M}$ is compact is a simplification and generally does not apply in deep learning. Local versions of Lemma 5 could be proved with $\mu,\alpha$ which are positive functions on the manifold, but not necessarily bounded away from zero. Naturally, this suffices in all cases where we provably remain in a local neighborhood in the course of optimization. We eschew this greater generality for the sake of geometric clarity and commit the pervasive sin of optimization theory for deep learning: We make global assumptions which can only be guaranteed locally. For a further comparison of geometric conditions in optimization and deep learning, see also Appendix C.
@@
Figure 3: Convexity analysis of $\phi(t)=L(w+tg)$ for $w$ near global minimizers of a loss function $L$ and $g=\nabla L(w)/\|\nabla L(w)\|$. Left: $\phi(t)$, middle: second derivative of $\phi$, right: estimated strong aiming parameter $\mu$ for $t\in[-1,1]$. Evidently, $\phi$ is strongly convex in a neighborhood of the minimizers. Strong aiming condition yields consistently larger constants than the strong convexity parameter obtained from second derivatives. Different colors correspond to different random initializations.
We illustrate in Figure 3 that our assumptions are locally reasonable in deep learning. We trained a fully connected neural network (with 10 layers, width 35, tanh activation) to fit labels $y_{i}$ at 100 randomly generated datapoints $x_{i}\in\mathbb{R}^{12}$. The small dataset size allowed us to use the exact gradient and loss function instead of stochastic approximations, for a better exploration of the loss landscape. Since the closest minimizer is generally unknown, we use the gradient as a proxy and examine the convexity of $\phi(t)=L(w+tg)$ for $w$ very close to the set of global minimizers of the loss function $L$ as in (2) and $g=\nabla L(w)/\|\nabla L(w)\|$. Labels were generated using a randomly initialized teacher network (with 7 layers and width 20). Student networks were trained for 10,000 epochs using stochastic gradient descent with Nesterov momentum, with learning rate $\eta=0.005$ and momentum $\rho=0.99$. Final training loss ranged between $10^{-12}$ and $10^{-9}$ across the five runs. Second derivatives were approximated using second order difference quotients $\phi^{\prime\prime}(t)\approx\frac{\phi(t+h)-2\phi(t)+\phi(t-h)}{h^{2}}$ for $h=0.01$. Similarly, the strong aiming parameter with respect to the global minimizer was estimated by $2\frac{\phi^{\prime}(t)t-\phi(t)+\inf\phi}{t^{2}}$ where $\phi^{\prime}(t)$ was estimated as $\frac{\phi(t+h)-\phi(t-h)}{2h}.$
Due to the inherent randomness of training, geometric behaviors varied. Generally, the loss function was convex along the gradient direction near minimizers, but neighborhood size, magnitude of the second derivative, and steepness of the loss function varied. Some runs exhibited convex but not strongly convex behavior, while others failed to reach zero loss along the line $w+tg/\|g\|$. The gradient is an imperfect approximation of the minimizer’s direction, and rounding errors may occur when it is so small near the set of minimizers.
3 Main Contributions
3.1 Optimization in continuous time